• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fatiha

    Log entry time 00:01:19 on May 3,2011

    Entry number 357891

    keyword=Shift Summary

    Shift crew: SL: F. Benmokhtar, TO: A. Daniel, 3rd: D. Day
    ----------


    16:00 Started shift with MCC trying to bring us back after beam
    recovery. We asked for 1 uA, raster on.
    16:35 MCC called, they are now able to give us beam.
    17:00 Beam back, got 120 uA, production on 3He.
    18:36 Target changed to Dummy, 5 uA
    19:15 HRS_L set to 2.905 GeV/c (Kin 3.1)
    19:35 Target changed to Optics, 5 uA
    19:59 Target moved to 3He, started production running @120uA
    20:25 MCC called, they are having issues with the pockel-cell
    Maximum current they can deliver is 115 uA for now.



    Run list:
    ---------
    Right arm is stuck, so we keep it at 2.845 GeV/c

    Kin 3.2b for left arm:
    ---------------------
    run #
    4073 spot++ check at 1uA
    4074 Production on 3He, beam current = 120 uA
    4075 Production on 3He, beam current = 120 uA
    4076 Dummy target @ 40 uA DT =~ 11%
    4077 BeO for Right arm, 5 uA

    Kin 3.1 for left arm:
    ---------------------
    P_L = 2.905 GeV/c (P_R stuck at 2.845)

    4078 BeO, 5 uA
    4079 Optics, 5 uA
    4080 Production on 3He, 120 uA 4M coda events, DT = 11%
    4081 Production on 3He, 115 uA
    4082 Production on 3He, 115 uA
    4083 Production on 3He, 115 uA
    4084 Production on 3He, 115 uA
    4085 Production on 3He, 115 uA
    4086 Production on 3He, 115 uA 5M coda events
    4087 Production on 3He, 115 uA 5M coda events
    4088 Production on 3He, 115 uA 5M coda events
    4089 goes to next shift

    Total number of left arm physics events recoreded during
    this shift is: 3.26 M